# GenIAS: Generator for Instantiating Anomalies in time Series

### v1

@@
Abstract.
A recent and promising approach for building time series anomaly detection (TSAD) models is to inject synthetic samples of anomalies within real data sets. The existing injection mechanisms have significant limitations — most of them rely on ad hoc, hand-crafted strategies which fail to capture the natural diversity of anomalous patterns, or are restricted to univariate time series settings. To address these challenges, we design a generative model for TSAD using a variational autoencoder, which is referred to as a Generator for Instantiating Anomalies in Time Series (GenIAS). GenIAS is designed to produce diverse and realistic synthetic anomalies for TSAD tasks. By employing a novel learned perturbation mechanism in the latent space and injecting the perturbed patterns in different segments of time series, GenIAS can generate anomalies with greater diversity and varying scales. Further, guided by a new triplet loss function, which uses a min-max margin and a new variance-scaling approach to further enforce the learning of compact normal patterns, GenIAS ensures that anomalies are distinct from normal samples while remaining realistic. The approach is effective for both univariate and multivariate time series. We demonstrate the diversity and realism of the generated anomalies. Our extensive experiments demonstrate that GenIAS — when integrated into a TSAD task — consistently outperforms seventeen traditional and deep anomaly detection models, thereby highlighting the potential of generative models for time series anomaly generation.
Anomaly generator, Time series, Anomaly detection
1. Introduction
@@
Unsupervised methods for Time Series Anomaly Detection (TSAD) often learn a model for boundary of normal data without access to labeled anomalies. This reliance on normal data makes models vulnerable to unseen or diverse anomaly patterns, which are common in time series settings. This is particularly true in complex multivariate scenarios, where anomalous patterns may vary significantly across different dimensions or even be correlated.
Anomaly injection techniques are an increasingly popular approach to address the limitations of solely relying on normal-boundary learning. Models like CARLA (Darban et al., 2025), CutAddPaste (Wang et al., 2024), NCAD (Carmona et al., 2022), and COUTA (Xu et al., 2024) have demonstrated the potential of synthetic anomaly injection in improving detection capabilities. However, these models face notable limitations. CARLA, CutAddPaste, and COUTA are restricted to predefined anomaly types, such as point, seasonal, trend, and shapelet anomalies. CutAddPaste further employs a rigid augmentation strategy (of “cut-add-paste”), which may not align well with realistic patterns in complex applications. NCAD, on the other hand, relies on prior knowledge about anomalies to generate realistic anomalies. However, it is limited by the anomalies present in the datasets and hence may not be able to generate diverse types of anomalies. Finally, all the existing anomaly injection models are susceptible to slight variations in the time series, as they generate anomalies in the raw time series space.
We address these gaps by introducing the Generator for Instantiating Anomalies in Time Series (GenIAS), a novel generative model designed to produce diverse and realistic synthetic anomalies for TSAD. In contrast to previous approaches, GenIAS operates in the latent space rather than the raw time series space. This makes it less susceptible to minor variations in the raw time series and allows it to learn more generalized normal patterns. While generative models mainly used to generate normal data, we are able to leverage the power of variational neural models to generate anomalies that align well with the temporal properties of time series, and with correlations between variables if the input time series is multivariate. GenIAS builds upon a hybrid Temporal Convolutional Network-Variational Autoencoder (TCN-VAE) architecture to move beyond rigid, handcrafted augmentation strategies and maps the complex properties of time series data into a multidimensional latent space. By employing a learned perturbation mechanism applied to the latent space variance and injecting the perturbed patterns into different segments of the time series using a patching mechanism, GenIAS generates anomalies with varying scales and enhanced diversity, resulting in more realistic anomaly patterns. We use the term realism to refer to realistic anomaly patterns in this paper. Additionally, we employ an adjusted triplet loss function, guided by a min-max margin to guarantee that the generated anomalies are both distinct from normal data and realistic at the same time. Further, we introduce a novel variance-scaling approach, which enforces the compactness of latent representations, making the generated anomalies more distinguishable from normal samples. We demonstrate the effectiveness of GenIAS when extended as a TSAD model. Our key contributions are as follows:
  • We propose GenIAS, which is a novel anomaly generator designed to seamlessly handle both univariate time series (UTS) and multivariate time series (MTS). By employing deep generative techniques, GenIAS produces anomalies that are both diverse and realistic.
  • We propose a novel loss function for our VAE-based anomaly generator, enforcing compact latent representations via a tailored prior variance. This improves the separation between normal and anomalous data in latent space, enhancing anomaly generation (see Figure 1). A theoretical analysis supports its effectiveness.
  • We demonstrate GenIAS’s ability to generate diverse and realistic anomalies through two newly introduced metrics. The evaluation is conducted on four state-of-the-art TSAD models leveraging anomaly injection, across nine widely used MTS and UTS datasets.
  • We demonstrate GenIAS’s broad applicability by integrating it into a state-of-the-art TSAD model and comparing it with 17 baselines. Extensive experiments on nine diverse UTS and MTS datasets confirm its effectiveness, with performance gains directly linked to the quality of GenIAS’s anomaly generation.
Figure 1. An illustration of our key idea: Variance-scaling enforces latent space compactness in the latent space, strengthening the separation between normal and generated anomalous samples by GenIAS, in latent space. The bottom figure, with $\sigma_{\text{prior}}$ = 0.5, visually shows this effect.
2. Related Work
@@
The field of time series anomaly detection has seen remarkable progress, and include traditional statistical methods as well as state-of-the-art deep learning approaches (Schmidl et al., 2022; Audibert et al., 2022; Zamanzadeh Darban et al., 2024). Early efforts included models like One-Class SVM (OC-SVM) (Schölkopf et al., 1999), LOF (Breunig et al., 2000), and Isolation Forest (Fei Tony Liu and Zhou, 2008); however, these methods faced challenges in handling high-dimensional and sequential data. Deep learning models like Donut (Xu et al., 2018), LSTM-VAE (Park et al., 2018) and OmniAnomaly (Su et al., 2019) introduced temporal dependency modeling and probabilistic approaches in TSAD.
@@
Representation-based methods, have pushed TSAD forward by focusing on meaningful embeddings. TS2Vec (Yue et al., 2022) applies self-supervised learning to capture multi-level semantic representations, while DCdetector (Yang et al., 2023) employs dual attention for permutation-invariant representations. Recent models such as TimesNet (Wu et al., 2023) and TranAD (Tuli et al., 2022a) model complex dependencies by leveraging attention mechanisms. Methods like MTAD-GAT (Zhao et al., 2020) utilize graph attention networks to model multivariate dependencies and THOC (Shen et al., 2020a) incorporates hierarchical modeling to capture temporal patterns at multiple granularities. Anomaly Transformer (Xu et al., 2021) leverages anomaly-sensitive features through attention mechanisms.
@@
OE (Hendrycks et al., 2019) integrates external data during training to expose the model to out-of-distribution instances, thereby making it more powerful. NCAD(Carmona et al., 2022) adapts OE for time series by infusing mixed contextual and random point anomalies to better separate normal and anomalous patterns. Anomaly injection and perturbation methods have emerged as crucial strategies for enhancing TSAD models. Approaches like CutAddPaste (Wang et al., 2024) and COUTA (Xu et al., 2024) simulate diverse anomaly patterns during training. These techniques have also been integrated into advanced models like CARLA (Darban et al., 2025) to create negative pairs for contrastive learning, enhancing generalization to unseen anomaly types.
3. GenIAS Method
3.1. Problem Statement
Let $\mathcal{D}$ represent a time series dataset consisting of $N$ time series windows, where each window $\mathbf{X}\in\mathbb{R}^{T\times D}$ is a time series window of length $T$ with $D$ dimension(s). Assume that samples $X$ in $\mathcal{D}$ is drawn from an unknown data distribution $p(\mathbf{X})$. In TSAD datasets, the training data is typically unlabeled, and it is assumed normal since anomalies are rare and have negligible impact. The objective is to use a generative mapping $\Gamma:\mathbf{X}\rightarrow\widetilde{\mathbf{X}}$ to transform a normal sample $\mathbf{X}$ into an anomalous sample $\widetilde{\mathbf{X}}\in\mathbb{R}^{T\times D}$, such that the anomalous distribution $p(\widetilde{\mathbf{X}})$ satisfies $p(\widetilde{\mathbf{X}})\neq p(\mathbf{X})$. The anomalies must preserve essential temporal properties in the normal data while exhibiting deviations that make them statistically separated.
The first challenge is to define a perturbation mechanism on the latent representation $\mathbf{z}\in\mathbb{R}^{L}$ with $L$ dimensions, parameterized by $p(\mathbf{z}\mid\mathbf{X})$, to produce a shifted latent distribution $p(\widetilde{\mathbf{z}})$. This ensures that the anomalies are distinguishable, satisfying $p(\widetilde{\mathbf{z}})\neq p(\mathbf{z})$ while retaining plausibility by remaining within the data manifold. Furthermore, by perturbing the latent representations, our model is less susceptible to the slight deviations (e.g., noise) in the raw time series space. The second challenge is to densify the normal latent distribution $p(\mathbf{z})$, enabling the mapping $\Gamma(\mathbf{X})=\hat{\mathbf{X}}$ to accurately reconstruct normal samples and establish a clear boundary for generating anomalies. Together, addressing these challenges guarantees that the generated anomalies achieve a balance between diversity and realism. This approach is designed to handle univariate ($D=1$) and multivariate ($D>1$) time series.
3.2. Anomaly Generation
GenIAS is a generative model designed to learn compact representations of normal time series data and generate realistic anomalies by strategically perturbing the latent space. The model comprises three main components: an encoder, a latent space representation, and a decoder. It is trained end-to-end with an objective function that balances reconstruction accuracy, latent space regularization, and anomaly generation quality.
3.2.1. Model Workflow and Perturbation Mechanism
The input to the model is a time series window $\mathbf{X}$ of shape $T\times D$. The input $\mathbf{X}$ is normalized to ensure consistency across features, improving model stability during training. The model architecture combines a TCN and a VAE to combine the strengths of both approaches. The VAE learns a latent representation of normal data and perturbs it in the latent space to enhance robustness to noisy time series and improve the learning of normal patterns. The TCN captures temporal dependencies within the time series more effectively. The encoder maps the input sample $\mathbf{X}$ into a latent representation $\mathbf{z}$ by capturing temporal dependencies through stacked convolutional layers. These layers, interleaved with activation functions and dropout, effectively model complex temporal patterns while reducing overfitting. The encoder outputs the parameters $\mu$ (mean) and $\sigma^{2}$ (variance), which define a Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ in the latent space, where $\mu$ and $\sigma$ are vectors of size $L$, the same as the size of latent vector $\mathbf{z}$.
The latent space representation forms the core of the model. For normal samples, latent vectors $\mathbf{z}$ are sampled as: $\mathbf{z}\sim\mathcal{N}(\mu,\sigma^{2})$. To generate anomalous representations $\tilde{\mathbf{z}}$, the standard deviation $\sigma$ is perturbed while keeping $\mu$ unchanged:
(1) $\tilde{\mathbf{z}}=\mu+\psi\cdot(\sigma\odot\epsilon),$
where $\psi$ is the learned perturbation scale, and $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is a noise vector. Perturbing $\sigma$ instead of $\mu$ ensures that the anomalies deviate in a way that preserves the central tendency of the normal data, thereby maintaining realism. Altering $\mu$ could disrupt this central tendency, leading to unrealistic deviations that do not align with the underlying data manifold.
The decoder reconstructs the input sample from $\mathbf{z}$ and $\tilde{\mathbf{z}}$. Using transpose convolutional layers upscales the latent representations back to the original input dimensions while preserving the temporal structure. The decoder outputs the following:
  • $\hat{\mathbf{X}}$: The reconstructed normal window, derived from $\mathbf{z}$.
  • $\tilde{\mathbf{X}}$: The generated anomalous window, derived from $\tilde{\mathbf{z}}$.
Figure 2 illustrates the overall workflow of the model, including the TCN-VAE architecture and the perturbation mechanism. The training process optimizes a combination of losses to ensure accurate reconstruction, realistic anomaly generation, and a compact latent space learning (see Algorithm 1 for details).
@@ 3.2.1. Model Workflow and Perturbation Mechanism
Figure 2. The overall architecture of GenIAS: The model comprises an encoder, a structured latent space, and a decoder. The encoder, built with a TCN-VAE, maps input time series windows into a latent representation by learning temporal dependencies. The latent space regularization ensures compact normal representations, while anomalies are generated by perturbing the variance of the learned distribution without altering the mean, preserving realistic deviations. The decoder reconstructs both normal and perturbed windows using transpose convolutional layers.
3.2.2. Reconstruction Loss
This portion of the overall loss is designed to accurately reconstruct normal input samples. It is defined using the Mean Squared Error (MSE) between an input sample $\mathbf{X}$ and its reconstruction $\hat{\mathbf{X}}$:
(2) $\mathcal{L}_{\text{recon}}=\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}\frac{1}{T\cdot D}\sum_{t=1}^{T}\sum_{d=1}^{D}\big{(}\mathbf{X}[t,d]-\hat{\mathbf{X}}[t,d]\big{)}^{2},$
This loss penalizes the discrepancy between $\mathbf{X}$ and $\hat{\mathbf{X}}$, ensuring that the model captures the key temporal and structural features of normal samples.
3.2.3. Perturbation Loss
The perturbation loss ensures that the generated anomalies $\tilde{\mathbf{X}}$ are both distinct from normal samples $\mathbf{X}$ and exhibit realistic deviations. This is achieved through two complementary terms: a triplet loss to enforce distinctiveness and a regularization term to maintain realism.
The first term, the triplet loss, encourages the original time series sample ${\mathbf{X}}$ to be closer to the reconstructed sample $\hat{\mathbf{X}}$ than the generated anomalous sample $\tilde{\mathbf{X}}$, ensuring that the anomalies are sufficiently distinct from original time series. By maximizing the margin between $d(\mathbf{X},\hat{\mathbf{X}})$ and $d(\mathbf{X},\tilde{\mathbf{X}})$, where $d(x,y)$ denotes the MSE distance, this term pushes the anomalous samples further away from the normal samples in the raw time series space. We also ensure that this margin is at least $\delta_{\text{min}}$, thus the generated anomalies be more distinct.
The second term (the regularization term) prevents the anomalous samples $\tilde{\mathbf{X}}$ from deviating excessively from the original samples, thereby preserving realism. It penalizes deviations where distance $d(\mathbf{X},\tilde{\mathbf{X}})$ exceeds a maximum allowable threshold $\delta_{\text{max}}$. This ensures that the generated anomalies remain plausible and do not degenerate into unstructured or unrealistic noise.
The perturbation loss is formulated as:
(3) $\begin{split}\mathcal{L}_{\text{perturb}}=&\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}\Big{[}\max\big{(}d(\mathbf{X},\hat{\mathbf{X}})-d(\mathbf{X},\tilde{\mathbf{X}})+\delta_{\text{min}},0\big{)}\Big{]}\\ &+\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}\max\big{(}d(\mathbf{X},\tilde{\mathbf{X}})-\delta_{\text{max}},0\big{)},\end{split}$
Together, these terms ensure that the generated anomalies achieve the dual objectives of being distinct yet realistic.
3.2.4. Zero-Perturbation Loss
The zero-perturbation loss focuses on dimensions in the input sample $\mathbf{X}$ that are entirely zero. It ensures that anomalies $\tilde{\mathbf{X}}$ introduce meaningful perturbations in these dimensions, encouraging deviations from the original sample where appropriate. The loss is defined as:
(4) $\mathcal{L}_{\text{zero-perturb}}=\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}(\frac{1}{T\cdot|\mathcal{Z}|}\sum_{t=1}^{T}\sum_{d\in\mathcal{Z}}\big{(}d(\mathbf{X},\tilde{\mathbf{X}})+1)^{-1},$
where $\mathcal{Z}$ is the set of zero-valued dimensions in $\mathbf{X}$. By penalizing dimensions that deviate unrealistically in regions of zero value, this loss encourages the model to generate plausible anomalies while preserving the unique characteristics of zero-valued dimensions.
3.2.5. Enhanced KL Divergence Loss

### v2

@@
Abstract
Synthetic anomaly injection is a recent and promising approach for time series anomaly detection (TSAD), but existing methods rely on ad hoc, hand-crafted strategies applied to raw time series that fail to capture diverse and complex anomalous patterns, particularly in multivariate settings. We propose a synthetic anomaly generation method named Generator for Instantiating Anomalies in Time Series (GenIAS), which generates realistic and diverse anomalies via a novel learnable perturbation in the latent space of a variational autoencoder. This enables abnormal patterns to be injected across different temporal segments at varying scales based on variational reparameterization. To generate anomalies that align with normal patterns while remaining distinguishable, we introduce a learning strategy that jointly learns the perturbation scale and compact latent representations via a tunable prior, which improves the distinguishability of generated anomalies, as supported by our theoretical analysis. Extensive experiments show that GenIAS produces more diverse and realistic anomalies, and that detection models trained with these anomalies outperform 17 baseline methods on 9 popular TSAD benchmarks.
Machine Learning, ICML
1 Introduction
@@
Unsupervised TSAD methods often learn the boundary of normal data for detection without labeled anomalies, making them vulnerable to diverse or subtle anomaly patterns. This is particularly true in complex multivariate scenarios, where anomalous patterns can vary significantly across dimensions or involve intricate correlations.
Anomaly injection techniques are gaining popularity for enhancing normal-boundary learning by addressing the lack of labels. Models like CARLA (Darban et al., 2025), CutAddPaste (Wang et al., 2024), NCAD (Carmona et al., 2022), and COUTA (Xu et al., 2024) show the promise of synthetic anomaly injection in improving detection, but each faces notable limitations. For instance, CARLA, CutAddPaste, and COUTA are limited to predefined anomaly types, such as point, seasonal, and shapelet anomalies. CutAddPaste employs a fixed “cut-add-paste” augmentation strategy, which may produce unrealistic patterns of time series. NCAD relies on prior knowledge of anomalies to generate realistic anomalies, but the diversity of its generated anomalies is constrained by the types of anomalies present in the training data. Moreover, most existing anomaly injection models operate directly in raw time series spaces, without modelling their normal patterns. This can result in unrealistic alignment between injected anomalies and the original time series components, producing anomalies that are out of context or entangled with normal patterns, and thus failing to provide anomaly discriminative information for TSAD training.
We address these gaps by introducing GenIAS, a novel generative model designed to produce realistic and diverse synthetic anomalies for time series data. In contrast to existing methods that operate directly on raw time series, GenIAS performs anomaly generation in the latent space using a hybrid Temporal Convolutional Network–Variational Autoencoder (TCN-VAE) architecture. Through variational reparameterization, this design encodes the temporal and inter-variable dependencies of normal patterns in multivariate time series within a continuous latent space, enabling fine-grained and controllable perturbations for generating diverse and realistic anomalies. This is achieved by our novel learnable perturbation mechanism in the learned latent space, which modulates latent variance to induce anomalies with varying degrees of deviation across different time series components.
To produce anomalies that align with the temporal properties of the original time series while remaining distinguishable from normal patterns, we introduce a joint training strategy that learns both the latent distribution of normal time series and the perturbation mechanism by enforcing two constraints. First, we propose a perturbation loss that encourages deviations of reasonable magnitude, avoiding perturbations that are either too trivial or excessively large. Second, we introduce a compact loss with a tunable variance prior initialized to a small value (e.g., <1<1) to induce compact latent representations in VAEs, as illustrated in Fig. 1. Our theoretical analysis shows that such compactness reduces overlap between the generated anomalies and normal patterns when applying the proposed learnable perturbation.
GenIAS injects anomalies through two stages, where a learned perturbation is first applied to the latent variance of normal time series representations to produce anomalous latent samples that are decoded into anomalous windows. To better reflect realistic TSAD scenarios, where anomalies are often localized rather than spanning all dimensions, a deviation-based patching strategy is then applied to selectively inject anomalous components from the anomalous window into the original window on a per-dimension basis, while preserving normal context. To evaluate anomaly generation quality, we introduce two complementary representation-based metrics that assess realism, defined as proximity to real anomalies in representation space, and diversity of the generated anomalies. GenIAS achieves a better balance in anomaly generation quality, and detection models trained with its generated anomalies consistently outperform 17 baselines across nine univariate and multivariate datasets.
Our key contributions are as follows:
Figure 1: Effect of compactness. A smaller prior $(\sigma_{\text{prior}}=0.5$ vs $\sigma_{\text{prior}}=1.0)$ variance produces compact normal latent representations and improved separation from generated anomalies
  • We propose a novel synthetic anomaly generation method, GenIAS, that operates in the latent space of variational autoencoders and generates diverse anomalies with varying degrees of deviation across different time series components, validating anomaly injection via controllable latent space perturbation as a promising strategy for improving unsupervised TSAD.
  • We introduce an optimization strategy for GenIAS that jointly learns the latent distribution of time series data and the perturbation mechanism. This ensures that injected deviations are informative and distinguishable while remaining coherent with the normal time series.
  • We propose two metrics for evaluating the diversity and realism for measuring anomaly generation, where GenIAS demonstrates better and more balanced results compared to the synthetic anomaly generation methods. We further show that GenIAS-enabled detection models demonstrate superior performance compared to 17 baseline methods, translating improved anomaly generation into stronger TSAD performance.
2 Related Work
@@
The field of TSAD has advanced significantly, spanning traditional statistical methods to SOTA deep learning approaches (Schmidl et al., 2022; Audibert et al., 2022; Zamanzadeh Darban et al., 2024). Early efforts included models like One-Class SVM (OC-SVM) (Schölkopf et al., 1999), LOF (Breunig et al., 2000), and Isolation Forest (Fei Tony Liu and Zhou, 2008), which struggled with handling high-dimensional and sequential data. In contrast, deep learning models like Donut (Xu et al., 2018), LSTM-VAE (Park et al., 2018) and OmniAnomaly (Su et al., 2019) introduced temporal dependency modeling, advancing TSAD performance.
@@
Representation-based methods have pushed TSAD forward by focusing on meaningful embeddings. TS2Vec (Yue et al., 2022) applies self-supervised learning to capture multi-level semantic features, while DCdetector (Yang et al., 2023) employs dual attention for permutation-invariant representations. Recent models such as TimesNet (Wu et al., 2023) and TranAD (Tuli et al., 2022b) leverage attention mechanisms to capture complex dependencies. Methods like MTAD-GAT (Zhao et al., 2020) utilize graph attention networks to model multivariate dependencies, and THOC (Shen et al., 2020a) incorporates hierarchical modelling to capture temporal patterns across multiple scales. Anomaly Transformer (Xu et al., 2021) leverages anomaly-sensitive features through attention mechanisms.
@@
OE (Hendrycks et al., 2019) improves robustness by introducing out-of-distribution examples during training. NCAD (Carmona et al., 2022) extends this idea to time series by injecting mixed contextual and random point anomalies to better separate normal and anomalous patterns. Anomaly injection and perturbation methods have become key strategies in TSAD, with methods like CutAddPaste (Wang et al., 2024) and COUTA (Xu et al., 2024) simulating diverse anomaly patterns during training. These techniques are also used in CARLA (Darban et al., 2025) to create negative pairs for contrastive learning, enhancing generalization to unseen anomalies.
@@
Figure 2: Overall architecture of GenIAS: The model comprises an encoder, a structured latent space, and a decoder. The encoder, based on a TCN-VAE, maps input time series windows into a latent representation by learning temporal dependencies. Latent space regularization encourages compact normal representations, and anomalies are generated by perturbing the variance of the learned distribution. The decoder reconstructs both normal and perturbed windows using transpose convolutional layers.
3 GenIAS Method
Problem Definition. Let $\mathcal{D}$ be a time series dataset consisting of $N$ windows, where each window $\mathbf{X}\in\mathbb{R}^{T\times D}$ has length $T$ and $D$ variables, drawn from an unknown data distribution $p(\mathbf{X})$. In typical TSAD settings, $\mathcal{D}$ is dominated by normal samples, while anomalies are rare. The objective of anomaly generation is to learn a mapping $\Gamma:\mathbf{X}\rightarrow\widetilde{\mathbf{X}}$ that transforms a normal window $\mathbf{X}$ into an anomalous window $\widetilde{\mathbf{X}}\in\mathbb{R}^{T\times D}$. The generated anomalies should preserve the local temporal context of $\mathbf{X}$ while introducing controlled deviations that make $\widetilde{\mathbf{X}}$ distinguishable from normal patterns and informative for training TSAD models.
For anomaly generation via latent space perturbation, we identify two key considerations for generating anomaly-discriminative samples. The first is to design a latent space perturbation mechanism for $\mathbf{z}\in\mathbb{R}^{L}$, parameterized by $p(\mathbf{z}\mid\mathbf{X})$, that injects controlled deviations into the latent representation, which serves as the input to the decoder for construct anomalous windows, to generate anomalies that satisfy the above properties. The second is to densify normal representations in latent space, which reduces overlap between normal samples and generated anomalies. This is applicable to both univariate and multivariate settings.
3.1 Anomaly Generation using GenIAS
GenIAS Overview. The core of GenIAS is a novel learnable perturbation mechanism operating on a variational latent space that is explicitly learned through a tailored training strategy, where distributions of normal time series data and the perturbation mechanism are jointly optimized. This joint optimization is motivated by the above design considerations, ensuring that the learned perturbations align well with the context of the normal time series distribution while being informative for anomaly generation.
Generation Workflow. As shown in Figure 2, GenIAS adopts an encoder-decoder architecture, where a Temporal Convolutional Network (TCN) is used as the backbone to extract complex temporal patterns and dependencies through stacked convolutional layers interleaved with nonlinear activations and dropout. A variational autoencoder is then used to model a continuous latent Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$, where $\mu$ and $\sigma^{2}$ are $L$-dimensional vectors capturing the central tendency (mean) and spread (variance) of normal patterns, respectively, providing a basis for controlled anomaly generation. A normal window can be generated by first sampling a latent vector $\mathbf{z}=\mu+(\sigma\odot\epsilon)$, which is then passed to the decoder to reconstruct the corresponding time series window.
Anomaly Injection Pipeline. GenIAS employs a two-stage pipeline for anomaly generation, consisting of anomalous window generation and deviation-based patching. Learnable perturbations are applied during latent sampling of a trained VAE to generate anomalous latent vectors (Eq.1), which are decoded into anomalous windows (detailed in Algo. 1).
While the reconstructed anomalous window captures coherent abnormal patterns, directly replacing the full window can be insufficient or overly aggressive for TSAD, as real anomalies are often localized. We therefore apply a deviation-based patching strategy that selectively replaces entire dimensions from the normal trajectory with their anomalous counterparts only when a non-trivial deviation is detected. The patching mechanism is detailed in Sec.3.3.
Deviation via Variance Perturbation. Although VAEs are commonly used for modelling and generating normal data, we exploit their variational latent distributions to generate anomalies by adjusting the variance around the latent mean to produce desired deviations. We propose perturbing the variance of the learned VAE instead of the mean so that the generated anomalies follow the temporal property and central tendency of the times series. To generate anomalous representations $\tilde{\mathbf{z}}$, the standard deviation $\sigma$ is perturbed while keeping $\mu$ fixed:
$\vskip-3.0pt\tilde{\mathbf{z}}=\mu+\psi\cdot(\sigma\odot\epsilon),\vskip-2.0pt$ (1)
where $\psi$ is a learned perturbation scale, and $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is a noise vector. We find that perturbing the latent mean can introduce large deviations, but it violates the central tendency, often exhibiting out of context temporal properties that are less informative for representing realistic anomalies.
3.2 GenIAS’s Joint Learning
We jointly learn the perturbation scale and the time series latent distribution to ensure compatibility. In addition to the reconstruction loss for encoding the context of time series (Eq. 5), we introduce two constraints to tighten the variance of the normal latent distribution and to regularize an effective perturbation scale. Specifically, we enforce a compact normal manifold through a compact loss (Eq. 2) to reduce overlap between generated anomalies and normal data, and calibrate the perturbation scale to produce context-aligned deviations using a perturbation loss (Eq. 3). Intuitively, effective synthetic anomalies should deviate from normal patterns without leaving the local context, since overly subtle deviations label near normal samples as anomalies and introduce label noise, while overly large deviations often become out of context and uninformative.